OpenStudy (anonymous):

Solve the following system of linear equations x+y+5z=13 5x+5y+26z=67 x=_____? y=_______? Z=________?

OpenStudy (anonymous):

5x+5y+26z=67 (x+y+5z=13)*-5 -5x-5y-25z=-65 add z=2 can you continue from there?